# ANN vs SNN: A case study for Neural Decoding in Implantable Brain-Machine Interfaces

### v1

@@ III-C Input Spike Processing
    This method is depicted in Figure 5(a). This method is used by ANN and LSTM models, where ANN uses $T_{W}=200$ms while LSTM uses $T_{W}=32$ms. Efficient implementation of such firing rate calculation with overlapping windows are shown in [30].
@@
    Similar to the summation method, the sub-window method uses information over the latest $T_{W}$ bin window. However, instead of summing all the spikes, it provides firing rate information at an even shorter time-scale (or with finer resolution) of $T_{W}/m$. Thus, the feature computed from the i-th probe itself becomes a vector $\overline{x_{i}}(t_{k})=[r_{i}^{1}(t_{k}),r_{i}^{2}(t_{k})...r_{i}^{m}(t_{k})]$ with $m$ components corresponding to firing rates in each of the $m$ sub-windows (duration of integration in Eq. 2 is reduced to $T_{W}/m$). The sub-window method is illustrated in Figure 5(b) and is used by the ANN_3D and SNN_3D models with $T_{W}=200$ms and $m=7$. The feature vector $\overline{x}(t_{k})$ for ANN_3D is defined according to Equation 4 as follows:
@@
    Streaming Method (used in SNN_Streaming): The streaming method, as the name suggests, processes the incoming spike data as a continuous stream as seen in Figure 5(c). In this case, $T_{W}=st=T_{GT}=4$ ms implying no overlap between consecutive windows. This allows for a direct interface between the probes and the model, without the need of adding additional compute cost to our network like the two methods mentioned before. The input feature vector $\overline{x}(t_{k})$ is given by the following equation:
@@ III-D Filter: Adding memory at output
Figure 6: Frequency distribution of ground truth and two model outputs from [17] for 2 sec of data from the fourth file in Table I shows higher frequency content in prediction. This indicates a strategy to choose filter cut off frequency.
@@
Most of the NN models (with the exception of LSTM and SNN_Streaming) introduced in Section III-B operate on a window or chunk of inputs; providing these windows in any order would result in the same prediction. However, in real life the motor output is a smooth signal with a continuous trajectory. To understand this, we plot in Fig. 6 the frequency content of ground truth trajectories of a sample 2-sec waveform and compare it with predicted trajectories of two models from [17]. It is clear that the predictions have much higher frequency content indicating ground truth trajectories are smoother. In signal processing, this can be rectified by using a filter, which amounts to adding a memory of the past output. Among the different possible filters, a Bessel filter is chosen because of its linear phase response which is good for maintaining arbitrary waveform shapes. Three different filtering methods are tested in this work. First, we use forward (Fwd) filtering, which can achieve real-time filtering, but cannot be zero-phase. On the contrary, bidirectional (Bid) filtering can effectively eliminate phase distortions, but it is generally applicable to offline filtering since the whole waveform is needed before processing begins. To achieve a compromise, block bidirectional filter with a sliding window is applied, such that only a latency penalty of half window size is applicable. In this work, a window size of $16$ was chosen to limit the latency to $32$ ms while the order of filters and their cutoff frequencies were varied between $1-4$ and $0.05-0.5$ respectively to find the optimum for each model.
III-E Metrics
@@
In order to evaluate the performance of the models comprehensively in terms of cost vs performance, three metrics are used: (1) number of operations, (2) memory footprint, and (3) accuracy. Three types of operations are considered for (1) – multiply, add and memory read (since the energy for memory access often dominates the energy for computations[31]). For most NNs, each synaptic operation comprises a multiply and add (MAC) while for SNN_Streaming, we only have additions (AC). The number of operations is used as proxy for power/energy in this work since the actual energy ratio between these three operations depends on bit-width, process node and memory size; more accurate energy evaluations will be the subject of future work. For (2), memory footprint is evaluated from model size where every parameter is stored using a 32-bit float number. For (3), R-Squared score is a commonly used metrics for regression tasks[15, 17, 14], which is defined by Equation 7:
@@
Another set of important metrics for NN hardware are throughput and latency. We have not considered them here since the considered NN models are small enough so that the total time taken for evaluating the prediction is dominated by the input data accumulation time shown in Section III-C. However, we do touch upon this point later in Section V.
III-F Training & Testing Details
@@
All models are trained for $50$ epochs using the SNNTorch framework, with a learning rate of $0.005$, a dropout rate between $0.3-0.5$, and an L2-regularization value between $0.005-0.2$. AdamW is chosen as the optimizer, Mean Squared Error (MSE) loss is determined as the loss function, and a learning scheduler (cosine annealing schedule) is used after every epoch. For ANN, ANN_3D, and SNN_3D, data is shuffled with batch size of $512$ in training. For SNN_3D, the membrane potential resets every batch, while reset occurs at the beginning of each reach for membrane potential in SNN_streaming and hidden states in LSTM. The distribution of reach dutrations show most reaches completed in less than $4$ sec while some reaches being much longer, presumably due to the NHP not attending to the task. Similar to [17], reaches that exceed $8$ seconds in length are removed to improve the training performance. Leaky Integrate-and-Fire Neuron is used in SNNs, where the threshold and $\beta$ are learned during training and Arctan is applied as a surrogate function. The membrane potential of neurons ceases to reset in the last layer to enable regression. The velocity predicted by SNNs is determined by scaling the membrane potential of neurons with a learnable constant parameter. For validation and testing, data is input to the models in chronological order, and reset mechanisms only occur at the beginning. Filters are employed exclusively during the inference process.
IV Results
@@
To comprehensively examine the capability of different models, we performed multiple experiments and evaluated models using the metrics mentioned in Section III-E. All the results except memory access are obtained from the neurobench harness[17] that does automated evaluation of the models; memory access is estimated based on theoretical equations of weight fetches based on experimentally observed sparsity multiplying the number of weights on a per layer basis. The findings are presented pictorially using two pareto plots, first comparing the accuracy versus operations trade-off and the second comparing accuracy versus memory footprint (e.g. see Figures 8 and 10). For the pareto plots shown in this section, the following colour scheme is used:
  • red markers are models using forward (Fwd) filtering
  • green markers are models using bidirectional (Bid) filtering
  • orange markers are models using block Bid filtering
  • markers with dark border are tiny variants of base models
A tabular summary of all the experiments performed for our base models can be found in Table III.
IV-A Model Size Search
As mentioned earlier, it was found that networks deeper than 3 layers performed poorly and hence deeper models were excluded from this study. The number of neurons in each of the two hidden layers was determined by searching within a certain range ($N_{0}=N_{ch}$ and $N_{3}=2$ are fixed). We used the ANN to do this search due to its simplest network structure and resultant fast training. Each model complexity is characterized by the number of neurons. The results obtained by varying $N_{1}$ and $N_{2}$ are shown in the Figure 7. The text shown in the figure represents the different network architectures ($N_{1}-N_{2}$ combinations) tested. As expected, the $R^{2}$ initially increases with increasing number of neurons but starts decreasing after the number of neurons reaches a certain value due to overfitting. The best trade-off between $R^{2}$ and complexity is determined by the networks lying on the pareto curve. Therefore, the two models with $N_{1}-N_{2}$ values of $32-48$ and $16-32$ were selected as the ‘base’ and ‘tiny’ variants respectively for ANN. Same variant sizes were near optimal for ANN_3D and SNN_3D (we do not show these tradeoff curves for brevity), while for SNN_Streaming, base and tiny variants represented $N_{1}-N_{2}$ values of $32-48$ and $16-48$ respectively.
@@ IV-A Model Size Search
Figure 7: Complexity versus $R^{2}$ for different ANN models tested. $N_{1}-N_{2}$ values of $16-32$ and $32-48$ were on the pareto curve and chosen as ‘tiny’ and ‘baseline’ variants.
Figure 8: Baseline Pareto plots showing trade-offs for different models: a) Compute cost vs. accuracy b) Memory footprint vs. accuracy
IV-B K-Fold Cross Validation
@@
It is important to verify that the result will not vary significantly regardless of how the data is split. Hence, K-fold cross-validation is used to test all six files for three models (ANN, ANN_3D and SNN_3D). We divided the data into five parts, randomly selecting four parts as training and the other part was divided into validation and testing. The means and standard deviations of $R^{2}$ for the 5-fold experiment are shown in Table II. Low variance of the results for all 3 cases implies using one-fold data split for our experiment is reasonable and will give dependable results. As a comparison, the results in Table III does show that without filter, the decoding accuracy for SNN_3D is the best and ANN is the worst with ANN_3D between the two. Hence, we just use the single data split in [17] described earlier for the rest of the results.
TABLE II: K-fold Cross Validation
@@
IV-C Baseline result and effect of filtering
The pareto plots showing the baseline comparison can be seen in Figure 8. The filter order and cutoff frequency are chosen separately for different filtering methods and different models. For ANN, ANN_3D, SNN_3D and SNN_Streaming, filter order and cutoff frequency are $4$ and $0.05$ for bidirectional filtering, while in block bid filtering, order is selected as $2$. For forward filtering, the parameters are $1$ and $0.15$ for order and cutoff frequency respectively. However, these are different in LSTM and set at $2$ and $0.07$ for order and cutoff for all kinds of filtering methods. These optimal parameters were selected based on results in the validation set by doing a grid search over filter orders $1-4$ and cut-off frequencies in the range $0.05-0.5$ as mentioned earlier in Section III-D.
In terms of the models that forms the pareto front of operations vs. accuracy (Fig. 8(a)), we observe all of them use bidirectional filtering to achieve their high accuracy. While this is impractical in real-time decoding, these results can be taken as a gold standard for the neurobench suite at this time since they represent the highest reported accuracy so far. Along the pareto front, the two SNN variants, SNN_3D and SNN_Streaming show the biggest difference in terms of operations required ($\approx 100$x) and accuracies ($\approx 0.022$), while LSTM and ANN_3D are at intermediate positions on the pareto front. In terms of memory usage (Fig. 8(b)), the pareto front is dominated by SNNs of the two types. The ANN_3D models have highest memory usage due to their input dimension being expanded by $m$ times to $m\times N_{ch}$–the weights in the first layer are dominant for memory footprint since $N_{0}>>N_{1},N_{2},N_{3}$.
If we ignore models with Bid filtering due to them not being applicable in real-time, the next pareto frontier is dominated by models with block BiD filtering. Both these results point to the extreme efficiency of combining NN models with traditional filtering for motor decoding. In that case, the pareto frontier for operations vs. accuracy consists of LSTM at the high end and SNN_streaming at the low end with ANN_3D in the middle. For the case of memory vs. accuracy, LSTM and SNN_streaming retain their positions at the top and bottom of the pareto curve, while SNN_3D is in the intermediate part. Figure 9 plots the actual trajectory of a ground truth reach waveform, a prediction from ANN_3D, and a filtered version. It can be seen how the filtered waveform is smoother and resembles the more natural motion of the primate’s finger.
Looking deeper at the effects of filtering, we see that Fwd filtering provided very little gains to most of the NN models. Using the block Bid filter provided maximum gains of $\approx 0.05$, $0.04$ and $0.03$ in $R^{2}$ for ANN_3d, SNN_3D and ANN models, while the gains were lower ($\approx 0.02$ or less) for LSTM and SNN_streaming models. This is intuitively understandable given that recurrent models have an inbuilt longer memory. Also, we see that tiny variant of ANN with block Bid filtering achieves similar accuracy of $\approx 0.61$ as the tiny or base SNN_streaming without filtering at similar memory and $\approx 8X$ computations. This confirms our initial hypothesis that adding memory to ANN models can indeed make their performance similar to SNNs. However, the performance of SNN_streaming also improves with the addition of the filter making this combination a great choice for decoding with very low computational and memory resources.
Figure 9: Predicted trajectory of X-velocity with and without filter showing the smoothness introduced by the filter making it similar to natural motion.
IV-D 80% vs 50% Training Split
To assess the performance of models when the training data increases, we increase the baseline training data from 50% to 80% as done in [28]. The results are listed in Table III and plotted in Figure 10. As expected, the $R^{2}$ of all models is generally higher by $0.03-0.04$ compared to the 50% baseline training data, which shows a high capacity for future improvement with more. Similar to Fig. 8, Fig. 10 also has models with Bid or block Bid filtering on the pareto curve. After the data increases, LSTM and ANN_3D show the greatest accuracy improvement with an increase of $0.04$ in $R^{2}$. In addition, LSTM becomes one of the frontiers in the Pareto plot for both memory footprint and operations displacing SNN_3D off the pareto curve in Fig. 10(a) for operations.

### v2

@@ 3.3 Feature Extraction by Input Spike Processing
    This method is depicted in Figure 5(a). This method is used by ANN and LSTM models, where ANN uses $T_{W}=200$ms while LSTM uses $T_{W}=32$ms. Both these bin window sizes were obtained after optimization using the training and validation data. Generally, shorter time windows are preferred to capture fine temporal structure of spike trains[54] and reduce the latency of response [55]. However, with short time windows, LSTMs (or other recurrent models) can retain a memory about long-term history through their state variables while ANNs cannot. Hence, the time window for ANN needs to be longer than that of LSTM to retain sufficient information. In terms of hardware realization, while this accumulation of spikes is straightforward for non-overlapping bin windows, cases with overlap would require repeated operations with overlapping data in naive implementations. Efficient implementation of such firing rate calculation with overlapping windows are shown in [18] using recursion.
@@
    Similar to the summation method, the sub-window method uses information over the latest $T_{W}$ bin window. However, instead of summing all the spikes, it provides firing rate information at an even shorter time-scale (or with finer resolution) of $T_{W}/m$. Thus, the feature computed from the i-th probe itself becomes a vector $\overline{x_{i}}(t_{k})=[r_{i}^{1}(t_{k}),r_{i}^{2}(t_{k})...r_{i}^{m}(t_{k})]$ with $m$ components corresponding to firing rates in each of the $m$ sub-windows (duration of integration in Equation 2 is reduced to $T_{W}/m$). The sub-window method is illustrated in Figure 5(b) and is used by the ANN_3D and SNN_3D models with $T_{W}=200$ms and $m=7$. The feature vector $\overline{x}(t_{k})$ for ANN_3D is defined according to Equation 4 as follows:
@@
    Streaming Method (used in SNN_Streaming): The streaming method, as the name suggests, processes the incoming spike data as a continuous stream as seen in Figure 5(c). The streaming method aims to achieve a better tradeoff between accuracy and computations compared to SNN_3D. In this case, $T_{W}=st=T_{GT}=4$ ms implying no overlap between consecutive windows. This allows for a direct interface between the probes and the model, without the need of adding additional compute cost to our network like the two methods mentioned before. The input feature vector $\overline{x}(t_{k})$ is given by the following equation:
@@ 3.4 Filters for SNN
Figure 6: Frequency distribution of ground truth and two model outputs from [29] for 2 sec of data from the fourth file in Table 1 shows higher frequency content in prediction. This indicates a strategy to estimate the filter cut off frequency.
@@
Most of the NN models (with the exception of LSTM and SNN_Streaming) introduced in Section 3.2 operate on a window or chunk of inputs; providing these windows in any order would result in the same prediction. However, in real life the motor output is a smooth signal with a continuous trajectory. To understand this, we plot in Figure 6 the frequency content of ground truth trajectories of a sample 2-sec waveform and compare it with predicted trajectories of two models from [29]. It is clear that the predictions have much higher frequency content indicating ground truth trajectories are smoother. In signal processing, this can be rectified by using a filter, which amounts to adding a memory of the past output.
Three types of filters are compared to further explore the effect of filtering– Bessel filter, Butterworth filter and Chebyshev filter. Bessel filters provide a linear phase response resulting in constant delay, but it requires higher filter order to achieve same attenuation at high frequency compared to others. Butterworth filter with no ripples in the passband provides a maximum flat response in the passband, which means the filter will introduce minimal variations to the desired signal amplitude. However, the Butterworth filter has a wide transition band, which makes its rolloff gentle, and hence, the Chebyshev filter becomes a third alternative.
In terms of digital filter implementation, three different techniques were tested in this work. First, we tried forward (Fwd) filtering, which can achieve real-time filtering, but cannot have zero phase shift. The improvements achieved with this method was marginal and we do not consider them in the following parts. On the contrary, bidirectional (Bid) filtering can effectively eliminate phase distortions, but it is generally applicable to offline filtering since the whole waveform is needed before processing begins. To achieve a compromise, block bidirectional filter with a sliding window is applied, such that only a latency penalty of half block size is applicable. We vary the block size between $16-80$, the order of filters between $2-8$ and their cutoff frequencies in the range of $0.05-0.5$ respectively to find the optimum for each model.
3.5 Metrics
@@
In order to evaluate the performance of the models comprehensively in terms of cost vs performance, three metrics are used: (1) number of operations, (2) memory footprint, and (3) accuracy. Three types of operations are considered for (1) – multiply, add and memory read (since the energy for memory access often dominates the energy for computations[56]). For most NNs, each synaptic operation comprises a multiply and add (MAC) since the neuron activations and weights are not binary. On the other hand, for SNNs, the synaptic operations only involves accumulations (AC) because of the binary neuron activation. Note that the operations mentioned in this work refer to the operations between synaptic connections as detailed in Section 4.4, while the operations within neurons that determine the membrane potential are excluded. The number of operations is used as proxy for power/energy in this work since the actual energy ratio between these three operations depends on bit-width, process node and memory size; more accurate energy evaluations will be the subject of future work. For (2), memory footprint is evaluated from model size where every parameter is stored using a 32-bit float number. For (3), Coefficient of determination ($R^{2}$) is a commonly used metrics for regression tasks[19, 29, 17], which is defined by Equation 7:
@@
Another set of important metrics for NN hardware are throughput and latency. We have not considered them here since the considered NN models are small enough so that the total time taken for evaluating the prediction is dominated by the input data accumulation time shown in Section 3.3 and delay due to filtering. However, we do touch upon this point later in Section 3.4.
3.6 Training & Testing Details
@@
All models are trained for $50$ epochs using the SNNTorch framework, with a learning rate of $0.005$, a dropout rate between $0.3-0.5$, and an L2-regularization value between $0.005-0.2$. AdamW is chosen as the optimizer, Mean Squared Error (MSE) loss is determined as the loss function, and a learning scheduler (cosine annealing schedule) is used after every epoch. For ANN, ANN_3D, and SNN_3D, data is shuffled with batch size of $512$ in training. For SNN_3D, the membrane potential resets every batch, while reset occurs at the beginning of each reach for membrane potential in SNN_Streaming and hidden states in LSTM. The distribution of reach durations show most reaches completed in less than $4$ sec while some reaches being much longer, presumably due to the NHP not attending to the task. Similar to [29], reaches that exceed $8$ seconds in length are removed to improve the training performance. Leaky Integrate-and-Fire Neuron is used in SNNs, where the threshold and $\beta$ are learned during training and Arctan is applied as a surrogate function [57]. The membrane potential of neurons ceases to reset in the last layer to enable regression. The velocity predicted by SNNs is determined by scaling the membrane potential of neurons with a learnable constant parameter. For validation and testing, data is input to the models in chronological order, and reset mechanisms only occur at the beginning. Filters are employed exclusively during the inference process.
4 Results
@@
To comprehensively examine the capability of different models, we performed multiple experiments and evaluated models using the metrics mentioned in Section 3.5. All the results except memory access are obtained from the neurobench harness[29] that does automated evaluation of the models; memory access is estimated based on theoretical equations of weight fetches based on experimentally observed sparsity multiplying the number of weights on a per layer basis. The findings are presented pictorially using two pareto plots, first comparing the accuracy versus operations trade-off and the second comparing accuracy versus memory footprint (e.g. see Figure 14 and Figure 16). A tabular summary of all the experiments performed for our base models can be found in Table 2. Table 2 also compares the results with other published work using statistical methods such as Steady State Kalman filter (SSKF), Unscented Kalman filter (UKF), recurrent Exponential-Family Harmonium (rEFH) etc.
Table 2: Baseline Performance and Comparison with Prior Works using high accuracy filter configuration. Best performing filter in a class of SNN models is highlighted in bold font. Incremental accuracy improvement by using $80\%$ data over $50\%$ data is shown in parenthesis.
Activation Model Size
Models Split Filters R2 Sparsity Computes (((( kB ))))
MACs ACs Memory
Access
ANN 50% No Filter [29] 0.5818 0.7514 4969.76 0 5,179.46 26.5234
Block Bid Filtering 0.6165 0.7514 4974.76 0 5184.46 26.5429
Bid Filtering 0.6168 0.7514 4974.76 0 5184.46 26.5429
[…]
Block Bid Filtering 0.6456 0.7417 5005.25 0 5210.65 26.5429
Bid Filtering 0.6461 0.7417 5005.25 0 5210.65 26.5429
ANN_3D 50% No Filter [29] 0.6013 0.7348 11507.07 0 11,555.31 134.5234
Block Bid Filtering 0.6646 0.7348 11512.07 0 11560.31 134.5429
Bid Filtering 0.656 0.7348 11512.07 0 11560.31 134.5429
[…]
Block Bid Filtering 0.6859 0.7324 11681.22 0 11649.82 134.5429
Bid Filtering 0.6887 0.7324 11681.22 0 11649.82 134.5429
SNN_3D 50% No Filter [29] 0.6219 0 32256 0 39,057.79 33.1992
Block Bid Filtering 0.6729 0 32261 0 39,062.79 33.2187
Bid Filtering 0.6687 0 32261 0 39,062.79 33.2187
[…]
SNN_Streaming 50% No Filter 0.6112 0.7453 0 971.26 1195.28 25.32
Block Bid Filtering 0.6449 0.7453 0 976.26 1,200.28 25.3395
Bid Filtering 0.6458 0.7453 0 976.26 1,200.28 25.3395
[…]
LSTM 50% No Filter 0.6508 0 22687.97 0 22913.27 90.95
Block Bid Filtering 0.6711 0 22692.97 0 22918.27 90.96
Bid Filtering 0.6683 0 22692.97 0 22918.27 90.96
[…]
Block Bid Filtering 0.7046 0 22692.97 0 22918.27 90.96
Bid Filtering 0.7051 0 22692.97 0 22918.27 90.96
SNN 2D[29] 50% - 0.5805 0.9976 0 413.52 1503 28.56
SNN2[30] 50% - 0.6292 - - 202 1815 30
ELM [17] 50% - 0.5546 - 217202 - 217344 1140.625
rEFH_dynamic [19] 320s - 0.6319 - 3167 - 13000 229376
[…]
@@
Figure 7: Complexity versus $R^{2}$ for different ANN models tested. $N_{1}-N_{2}$ values of $16-32$ and $32-48$ were on the pareto curve and chosen as ‘tiny’ and ‘baseline’ variants. Here, model complexity is characterized by the number of synaptic weights.
4.1 Model Size Search
As mentioned earlier, it was found that networks deeper than 3 layers performed poorly and hence deeper models were excluded from this study. The number of neurons in each of the two hidden layers was determined by searching within a certain range ($N_{0}=N_{ch}$ and $N_{3}=2$ are fixed). We used the ANN to do this search due to its simplest network structure and resultant fast training. The results obtained by varying $N_{1}$ and $N_{2}$ are shown in the Figure 7. Here, model complexity is characterized by the number of synaptic weights. The text shown in the figure represents the different network architectures ($N_{1}-N_{2}$ combinations) tested. As expected, the $R^{2}$ initially increases with increasing number of neurons but starts decreasing after the number of neurons reaches a certain value due to overfitting. The best trade-off between $R^{2}$ and complexity is determined by the networks lying on the pareto curve shown in blue in Figure 7. Therefore, the two models with $N_{1}-N_{2}$ values of $32-48$ and $16-32$ were selected as the ‘base’ and ‘tiny’ variants respectively for ANN. Same variant sizes were near optimal for ANN_3D and SNN_3D (we do not show these tradeoff curves for brevity), while for SNN_Streaming, base and tiny variants represented $N_{1}-N_{2}$ values of $32-48$ and $16-48$ respectively.
4.2 K-Fold Cross Validation
@@
It is important to verify that the result will not vary significantly regardless of how the data is split. Hence, K-fold cross-validation is used to test all six files for three models (ANN, ANN_3D and SNN_3D). We divided the data into five parts, randomly selecting four parts as training and the other part was divided into validation and testing. The means and standard deviations of $R^{2}$ for the 5-fold experiment are shown in Table 3. Low variance of the results for all 3 cases implies using one-fold data split for our experiment is reasonable and will give dependable results. As a comparison, the results in Table 2 does show that without filter, the decoding accuracy for SNN_3D is the best and ANN is the worst with ANN_3D between the two. Hence, we just use the single data split in [29] described earlier for the rest of the results.
Table 3: K-fold Cross Validation
@@
SNN_Streaming 0.6144 0.027
LSTM 0.6755 0.036
4.3 Filtering: Performance improvement and Optimization
Figure 8: Comparison of different $4$-th order filters using bidirectional technique and cut-off frequency of $0.05$. Bessel filters show the best improvement. Statistical significance tested using paired t-test.
Figure 9: Comparison of three types of filters in terms of velocity and position based on one reach in the file "$indy\_20160622\_01"$. The three columns indicate SNN with Bessel, Butter, and Chebyshev filter respectively. The first row shows the X velocity, the second row shows the Y velocity. 4-th order filter with block size of 32 and cut-off frequency of 0.05 is used.
Figure 10: Comparison of block size versus $R^{2}$ in different filter order: a) The performance for SNN_3D using bessel filter with cutoff frequency of 0.05. b) The performance for SNN_Streaming using bessel filter with cutoff frequency of 0.05.
Figure 11: Latency versus $R^{2}$ for block Bid filtering using $2$-nd order Bessel filters with cut-off frequency of $0.05$.The number denotes different block sizes. Block size of $32$ gives an optimal tradeoff between accuracy and latency.
First, we compare the performance of different types of filters; results for bidirectional filtering are shown here but similar conclusion holds for the Block bidirectional case as well. The comparison of these three filters with the original result is depicted in Figure 8. Here, the order of the filter is fixed at $4$ and the cut-off frequency is $f_{c}=0.05$ based on the estimate from Figure 6. It is clear that the Bessel filter provides the maximum improvement of $R^{2}$ both in SNN_3D and SNN_Streaming model (statistical significance tested using paired t-test), which is about 8.2% higher than SNN_3D and 5.7% higher than SNN_Streaming without using filters. The constant delay property of Bessel filters is crucial in not distorting the waveform shape. To visualize this qualitatively, X and Y velocities for one reach from one file after filtering by the three filters is shown in Figure 9(reaches from other files are shown in the Supplementary material). It can be seen that the velocity trajectories after Chebyshev filtering is not as smooth as the other two with Bessel filters producing the smoothest and most natural trajectories. Thus, we choose Bessel filters to perform the rest of the experiments in this paper.
To find the combination of filter order and block size with the best trade-off in terms of accuracy and latency, we evaluated the performance for SNN_3D and SNN_Streaming using block Bid filtering of bessel filter with the cut-off frequency of $0.05$. The block size is varied between $16-80$, while the filter order is swept between $2-8$. The results for both types of SNN models shown in Figure 10 indicate a large increase in $R^{2}$ when the block size increases to $32$ and marginal improvements from there on, making block size of $32$ a good choice. For block sizes of $32$, a filter order of $4$ is optimal (statistically significant difference between order 2 and order 4 tested at the $5\%$ level using paired t-test for both SNN_3D and SNN_Streaming), while for higher block sizes, filter order of $6$ gives better results. There are no results for block size of $16$ with filter order larger than $4$ because of insufficient samples to perform the filtering. In this case, filter order of $2$ gives the best accuracy (statistically significant difference between order 2 and order 4 exists using the paired t-test at the $5\%$ level for SNN_Streaming and $10\%$ level for SNN_3D). From these results, a block size of $32$ and a filter order of $4$ seems like the best choice. However, there is a direct tradeoff between filter order and latency of the output as discussed next; this may make block size of $16$ more appropriate for some applications[30].
Latency between input and output is important for real-time applications with closed-loop operation such as motor decoding. The total time, $T_{tot}$, taken to produce an output by a NN decoder is given by $T_{tot}=St+T_{comp}$ where $St$ is the stride to capture the new input data ($=T_{GT}=4$ ms in this work) and $T_{comp}$ is the time taken to process the computations in the neural network. Given the very fast and energy-efficient In-memory computing (IMC) approaches to implement NN models prevalent now[58, 59] and the small networks considered in this paper, we can assume $T_{comp}<<St$ making the throughput almost entirely dependent on $St$, i.e. time taken to capture new neural input spikes. Note that the bin window, $T_{W}$ does not add any extra penalty on latency of output generation; however, after every change of target, the prediction will be inaccurate for a time related to $T_{W}$ to allow enough relevant input to fill up the bin window. However, output filtering may induce an extra penalty on the latency. Bid filters produced best results as seen in the Table 2; however, they cannot be employed in real-time applications since they need to store the raw data in memory first and then apply forward filtering two times in opposite directions. The block Bid filter is chosen as a compromise where the filter window is used to determine the length or block of samples that are filtered at one time, and the predicted point is located at the center of the sample window. Thus, the latency introduced by the block Bid filter is theoretically equal to half the length of the filter window. Figure 11 compares the accuracy of 2nd order Bessel filters for various block sizes. This shows that the minimum latency achievable by a 2nd order Bessel filter with block size $16$ is $32$ ms for this dataset with $T_{GT}=4$ ms. Hence, we select two Bessel filter configurations for further simulations–lower latency (2nd order, block size $16$) and higher accuracy (4th order, block size $32$).
Lastly, we have so far chosen the cutoff frequency $f_{c}$ in an adhoc fashion based on Figure 6. However, $f_{c}$ also can be optimized for the above two configurations. SNN models are not directly involved in the training process; rather, the input data is processed by pre-trained SNNs and the model output is then passed to the Bessel filter. Two approaches were tried–one where the digital filter is treated as a single layer neural network and another where $f_{c}$ is directly optimized by using search algorithms, with the latter providing better results. Grid search is used first to determine a reasonably accurate range around $0.05$ for $f_{c}$ with the potential maximum of $R^{2}$. Then search step and range are halved, based on the preferred range found by grid search. This process is repeated independently for each of the six files until $R^{2}$ stops increasing. Figure 12 shows the optimization results for both SNN_Streaming and SNN_3D for the two configurations. The results indicate a statistically significant difference between filtered and unfiltered data exists using the paired t-test at the 5% level for SNN_3D and 10% level for SNN_Streaming.
Figure 12: Cut-off frequency optimization for the two cases: one case of filter order of 2 and block size of 16, another case of filter order of 4 with block size of 32, respectively. Statistical significance tested using paired t-test.
4.4 Baseline result and Pareto plots
Figure 13: Comparison of baseline performance across various models, both using block bid filtering and without filtering. The statistically significant difference between filtered and unfiltered data exists using the paired t-test at the 5% level for ANN and SNN models.